# Closed-Form Evaluation of $\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})$ via Infinite Products

### v3

@@ 1 Introduction
and the decimal expansion of $h(2)=\tfrac{1}{2}\ln(\sinh(\pi)/\pi)$ appears in the OEIS as A372862. We shall prove (Theorem 27) that $h(3)=\frac{1}{2}\ln(3/2)$ and that this is the unique value of $k$ for which $h(k)$ is the logarithm of a rational number.
@@ 1.1 Main results
$\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right)=\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right),$ (2)
@@
$f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right)\quad\text{and}\quad g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right)$ (3)
are given by explicit formulas in terms of the gamma function (for odd $k$) or trigonometric functions (for even $k$); see Theorem 5.
@@
which follows from the Taylor expansion of $\operatorname{arctanh}$ and the Fubini–Tonelli theorem (see Remark 9).
@@ 1.3 Structure of the paper
Section 2 recalls the infinite product formulas and establishes preliminary results. Section 3 contains the proof of Theorem 1. Section 4 derives the zeta function identity and studies the correction function $C(k)$. Section 5 establishes the representation for $\gamma$ using Frullani’s theorem. Section 6 develops the analytic properties of $h(k)$, including monotonicity, convexity, asymptotic behavior, and two-sided bounds. Section 7 presents an exponentially convergent reformulation of the $\gamma$ representation and establishes tail bounds for the direct computation of $h(k)$. Section 8 proves the transcendence of $h(3)$ and discusses the arithmetic nature of $h(k)$ for other $k$. Section 9 collects numerical examples.
@@ 2 Infinite product formulas
$f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right)\quad\text{and}\quad g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right).$ (8)
@@
The decimal expansion of this constant appears in the OEIS as A372862.
To connect these products to our desired sums, we require the following lemmas relating logarithms of products to series involving the Riemann zeta function.
@@ 3 Proof of the main theorem
Subtracting the identity in Lemma 8 from that in Lemma 7 gives
$\ln\!\left(\frac{f(k)}{g(k)}\right)=\sum_{n=1}^{\infty}\frac{\bigl[(-1)^{n+1}+1\bigr]\zeta(kn)-1}{n}.$
The coefficient $(-1)^{n+1}+1$ equals $2$ when $n$ is odd and 0 when $n$ is even. Therefore
$\ln\!\left(\frac{f(k)}{g(k)}\right)=\sum_{n=1}^{\infty}\frac{2\bigl[\zeta\bigl(k(2n-1)\bigr)-1\bigr]}{2n-1}+\sum_{n=1}^{\infty}\frac{1}{2n-1}.$ (15)
Because the second sum is divergent, we regroup in adjacent odd-even blocks without reordering terms to keep all sums convergent:
$\ln\!\left(\frac{f(k)}{g(k)}\right)=2\zeta(k)-1-\sum_{n=1}^{\infty}\left[\frac{1}{2n}+\frac{1-2\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}\right].$
Now, for the sum portion,
$\sum_{n=1}^{\infty}\left[\frac{1}{2n}+\frac{1-2\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}\right]$ $=\sum_{n=1}^{\infty}\left[\frac{2n+1}{(2n)(2n+1)}+\frac{2n\!\left(1-2\zeta\!\bigl(k(2n+1)\bigr)\right)}{(2n)(2n+1)}\right]$
$=\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)}+2\sum_{n=1}^{\infty}\frac{1-\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}.$
Also,
$\sum_{n=1}^{\infty}\frac{1}{2n(2n+1)}=\sum_{n=1}^{\infty}\left(\frac{1}{2n}-\frac{1}{2n+1}\right)=1-\ln 2.$
Hence
$\ln\!\left(\frac{f(k)}{g(k)}\right)=2\zeta(k)-2+\ln 2-2\sum_{n=1}^{\infty}\frac{1-\zeta\!\bigl(k(2n+1)\bigr)}{2n+1}.$ (16)
For $s>1$, we have $1-\zeta(s)=-\sum_{m=2}^{\infty}m^{-s}$. By the Fubini–Tonelli theorem,
$\sum_{n=1}^{\infty}\frac{1-\zeta\bigl(k(2n+1)\bigr)}{2n+1}=-\sum_{m=2}^{\infty}\sum_{n=1}^{\infty}\frac{1}{(2n+1)\,m^{k(2n+1)}}.$ (17)
The key observation is that for $|x|<1$, $\operatorname{arctanh}(x)=\sum_{n=0}^{\infty}x^{2n+1}/(2n+1)$, so
$\sum_{n=1}^{\infty}\frac{(m^{-k})^{2n+1}}{2n+1}=\operatorname{arctanh}(m^{-k})-m^{-k}.$
Substituting into (17) and then into (16) gives
$\ln\!\left(\frac{f(k)}{g(k)}\right)$ $=2\zeta(k)-2+\ln 2+2\sum_{m=2}^{\infty}\bigl[\operatorname{arctanh}(m^{-k})-m^{-k}\bigr]$
$=2\zeta(k)-2+\ln 2+2\sum_{m=2}^{\infty}\operatorname{arctanh}(m^{-k})-2\bigl[\zeta(k)-1\bigr]$
$=\ln 2+2\sum_{m=2}^{\infty}\operatorname{arctanh}(m^{-k}).$
Therefore
$\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right)=\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right),$
which yields the result. ∎
@@
For every integer $k\geq 2$, the sum $h(k)$ admits the equivalent representations
$h(k)$ $=\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right)$ (18)
@@
Here (19) is Theorem 1, while (20) follows from the Taylor expansion $\operatorname{arctanh}(x)=\sum_{m=0}^{\infty}x^{2m+1}/(2m+1)$ applied with $x=n^{-k}$ and exchange of summation (justified by absolute convergence for $k\geq 2$).
@@ 4 Connection to the Riemann zeta function
From the proof of Theorem 1, we obtained
$\frac{1}{2}\ln\!\left(\frac{f(k)}{g(k)}\right)=\zeta(k)-1+\sum_{n=1}^{\infty}\frac{\zeta\bigl(k(2n+1)\bigr)-1}{2n+1}+1-\ln 2+\frac{1}{2}.$
More directly, combining (20) with the definition of $C(k)$ gives
$h(k)=\bigl[\zeta(k)-1\bigr]+C(k),$
@@ 5.3 Proof of the gamma formula
Integrating $C^{\prime}(k)$ from 11 to $k$ using Lemmas 13 and 14:
@@ 8.3 Uniqueness of $k=3$
The value $k=3$ is the unique integer $k\geq 2$ for which rationality follows from the cyclotomic telescoping mechanism.
@@
We treat even and odd $k$ separately.
Even $k$. For even $k\geq 2$, the product formulas of Theorem 5 express $f(k)$ and $g(k)$ in terms of $\sin$ and $\sinh$ evaluated at algebraic multiples of $\pi$. In particular, for $k=2$ we have $f(2)/(2g(2))=\sinh(\pi)/\pi=(e^{\pi}-e^{-\pi})/(2\pi)$. By Nesterenko’s theorem [5], $\pi$ and $e^{\pi}$ are algebraically independent over $\mathbb{Q}$, so any nonzero element of the field $\mathbb{Q}(\pi,e^{\pi})$ is transcendental. Since $\sinh(\pi)/\pi\in\mathbb{Q}(\pi,e^{\pi})\setminus\{0\}$, the quantity $f(2)/(2g(2))$ is transcendental, and in particular irrational. The argument for general even $k$ is analogous: the product $f(k)/(2g(k))$ lies in a field generated over $\mathbb{Q}$ by $\pi$ and exponentials $e^{\pi\alpha}$ for algebraic $\alpha$, and is expected to be transcendental; at minimum, it is not known to be rational in these cases.
Odd $k$, $k=3$. This is the content of Theorem 27.
Odd $k\geq 5$. The telescoping that occurs for $k=3$ depends on the cyclotomic identity $\Phi_{6}(n+1)=\Phi_{3}(n)$, which is special to the quadratic case: $(n+1)^{2}-(n+1)+1=n^{2}+n+1$. For odd primes $p\geq 5$, the polynomials $\Phi_{p}$ and $\Phi_{2p}$ both have degree $\varphi(p)=p-1\geq 4$, and $\Phi_{2p}(n+1)\neq\Phi_{p}(n)$; for example, the coefficient of $n^{p-2}$ in $\Phi_{2p}(n+1)$ is $\binom{p-1}{1}-1=p-2$, whereas in $\Phi_{p}(n)$ it is $1$. Hence
@@
Theorem 30 shows that $f(k)/(2g(k))$ is transcendental for even $k$ and irrational for $k=3$. However, the irrationality of $h(k)=\frac{1}{2}\ln\bigl(f(k)/(2g(k))\bigr)$ for even $k$ remains open: since $f(k)/(2g(k))$ is transcendental in this case, Baker’s theorem does not apply, and proving that $\ln(\alpha)$ is irrational for a specific transcendental number $\alpha$ appears to require algebraic independence results beyond those currently available. For odd $k\geq 5$, even the algebraic nature of $f(k)/(2g(k))$ is unresolved; should this quantity prove to be algebraic, Baker’s theorem would immediately yield the transcendence of $h(k)$.
@@ 9.5 OEIS sequence data
The decimal expansion of $h(2)=\frac{1}{2}\ln(\sinh(\pi)/\pi)$ appears in the OEIS as A372862. The Euler–Mascheroni constant $\gamma$ is A001620. The values of $\zeta(k)$ for small $k$ correspond to well-known constants: $\zeta(2)=\pi^{2}/6$ (A013661), $\zeta(3)$ is Apéry’s constant (A002117), $\zeta(4)=\pi^{4}/90$ (A013662), and so on.

### v4

@@ 1 Introduction
and the decimal expansion of $\ln(\sinh(\pi)/\pi)=2h(2)$ appears in the OEIS as A393668. We shall prove (Theorem 29) that $h(3)=\frac{1}{2}\ln(3/2)$ and that this is the unique value of $k$ for which $h(k)$ is the logarithm of a rational number.
@@ 1.1 Main results
$\sum_{n=2}^{\infty}\operatorname{arctanh}\!\left(\frac{1}{n^{k}}\right)=\frac{1}{2}\ln\!\left(\frac{g(2k)}{g(k)^{2}}\right),$ (2)
@@
$g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right)$ (3)
admits explicit formulas in terms of the gamma function (for odd $k$) or trigonometric functions (for even $k$); see Theorem 5. Equivalently, defining
$f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right),$ (4)
we have
$h(k)=\frac{1}{2}\ln\!\left(\frac{f(k)}{2\,g(k)}\right),$ (5)
since $f(k)=2g(2k)/g(k)$.
The formulation (2) in terms of $g$ alone is the more natural one: it shows that $h(k)$ measures the deviation of $g$ from a doubling law $g(2k)=g(k)^{2}$. Equivalently,
$g(2k)=g(k)^{2}\,e^{2h(k)}.$ (6)
@@
which follows from the Taylor expansion of $\operatorname{arctanh}$ and the Fubini–Tonelli theorem (Proposition 9).
@@ 1.2 Related work
1.3 Novelty and imported tools
The product formulas for $f(k)$ and $g(k)$ (Theorem 5) are due to Cantrell and rest on classical Weierstrass–Euler–Gauss theory for $\Gamma$; these are imported and provide the starting point. The Baker and Nesterenko transcendence theorems are applied as black boxes. The main contribution of this paper is Theorem 1, which connects Cantrell’s product identities to arctanh sums, yielding the closed-form evaluation $h(k)=\frac{1}{2}\ln\bigl(g(2k)/g(k)^{2}\bigr)$. Further new results include the Euler–Mascheroni representation and its accelerated variant (Corollary 3, Theorem 25).
@@ 1.4 Structure of the paper
Section 2 recalls the infinite product formulas and establishes preliminary results. Section 3 contains the proof of Theorem 1 and the zeta-series representation (Proposition 9). Section 4 introduces the companion series and derives related identities. Section 5 derives the zeta function identity and studies the correction function $C(k)$. Section 6 establishes the representation for $\gamma$ using Frullani’s theorem. Section 7 develops the analytic properties of $h(k)$, including monotonicity, convexity, asymptotic behavior, and two-sided bounds. Section 8 presents an exponentially convergent reformulation of the $\gamma$ representation and establishes tail bounds for the direct computation of $h(k)$. Section 9 proves the transcendence of $h(3)$ and discusses the arithmetic nature of $h(k)$ for other $k$. Section 10 collects numerical examples.
@@ 2 Infinite product formulas
$g(k):=\prod_{n=2}^{\infty}\!\left(1-\frac{1}{n^{k}}\right)\quad\text{and}\quad f(k):=\prod_{n=1}^{\infty}\!\left(1+\frac{1}{n^{k}}\right).$ (11)
@@
The decimal expansion of $\ln(\sinh(\pi)/\pi)=2h(2)$ appears in the OEIS as A393668.
The following lemmas express the logarithms of $f(k)$ and $g(k)$ as zeta series; they are used in the proof of Theorem 11 and other consequences.
@@ 3 Proof of the main theorem
The argument is direct: express each arctanh term as a log-ratio, sum the logarithms (justified by absolute convergence of the underlying product), and identify the result with $f(k)/(2g(k))$.
@@
For $n\geq 2$ and $k\geq 2$, we have $0<n^{-k}<1$, so
$\operatorname{arctanh}(n^{-k})=\frac{1}{2}\ln\frac{1+n^{-k}}{1-n^{-k}}.$
Since $\sum_{n=2}^{\infty}|\ln(1\pm n^{-k})|<\infty$ for $k\geq 2$ (by comparison with $\sum n^{-k}$), we may interchange summation with the logarithm:
$\sum_{n=2}^{\infty}\operatorname{arctanh}(n^{-k})=\frac{1}{2}\sum_{n=2}^{\infty}\ln\frac{1+n^{-k}}{1-n^{-k}}=\frac{1}{2}\ln\prod_{n=2}^{\infty}\frac{1+n^{-k}}{1-n^{-k}}.$
From the definitions, $f(k)=(1+1)\prod_{n=2}^{\infty}(1+n^{-k})=2\prod_{n=2}^{\infty}(1+n^{-k})$ and $g(k)=\prod_{n=2}^{\infty}(1-n^{-k})$, so
$\prod_{n=2}^{\infty}\frac{1+n^{-k}}{1-n^{-k}}=\frac{f(k)/2}{g(k)}=\frac{f(k)}{2\,g(k)},$
which gives (5). Finally, using $1-n^{-2k}=(1-n^{-k})(1+n^{-k})$,
$g(2k)=\prod_{n=2}^{\infty}(1-n^{-2k})=\prod_{n=2}^{\infty}(1-n^{-k})(1+n^{-k})=g(k)\cdot\frac{f(k)}{2},$
so $f(k)/(2g(k))=g(2k)/g(k)^{2}$, and (2) follows. ∎
Proposition 9 (Zeta-series representation).
For all integers $k\geq 2$,
$h(k)=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}.$ (18)
The series converges absolutely with tail bound $O(2^{-3k})$.
Proof.
Expand $\operatorname{arctanh}(n^{-k})=\sum_{m=0}^{\infty}n^{-k(2m+1)}/(2m+1)$ for each $n\geq 2$. The double sum $\sum_{n\geq 2}\sum_{m\geq 0}n^{-k(2m+1)}/(2m+1)$ converges absolutely for $k\geq 2$ (dominated by $\sum_{n,m}n^{-2(2m+1)}$), so the Fubini–Tonelli theorem permits interchange:
$h(k)=\sum_{m=0}^{\infty}\frac{1}{2m+1}\sum_{n=2}^{\infty}\frac{1}{n^{k(2m+1)}}=\sum_{m=0}^{\infty}\frac{\zeta\bigl((2m+1)k\bigr)-1}{2m+1}.\qed$
@@
Theorem 1 and Proposition 9 together give three equivalent closed forms for $h(k)$:
$h(k)$ $=\frac{1}{2}\ln\!\left(\frac{g(2k)}{g(k)^{2}}\right)$ (19)
@@
4 The companion series
The zeta-series representation (18) sums over the odd-indexed terms $\zeta((2m+1)k)$. There are three natural companions: the all-index sum, the even-index sum, and the alternating sum. All admit closed forms in terms of $g$.
Theorem 11 (Companion series).
For $k\geq 2$, define
$A(k)$ $:=\sum_{n=1}^{\infty}\frac{\zeta(nk)-1}{n},$ (22)
$E(k)$ $:=\sum_{n=1}^{\infty}\frac{\zeta(2nk)-1}{2n},$ (23)
$B(k)$ $:=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}\bigl[\zeta(nk)-1\bigr]}{n}.$ (24)
Then
$A(k)$ $=-\ln g(k),$ (25)
$E(k)$ $=-\tfrac{1}{2}\ln g(2k),$ (26)
$B(k)$ $=\ln\bigl(f(k)/2\bigr).$ (27)
Moreover, $h(k)$ and the three companions satisfy
$A=h+E,\quad B=h-E,\quad h=\tfrac{1}{2}(A+B),\quad E=\tfrac{1}{2}(A-B).$ (28)
Proof.
Equation (25) is Lemma 8. For (26), substitute $2k$ for $k$ in Lemma 8: $A(2k)=\sum_{n\geq 1}(\zeta(2nk)-1)/n=-\ln g(2k)$. Since $E(k)=\frac{1}{2}A(2k)$, we get $E(k)=-\frac{1}{2}\ln g(2k)$.
For (27), write $\ln(f(k)/2)=\ln f(k)-\ln 2$. From Lemma 7,
$\ln f(k)=\sum_{n=1}^{\infty}\frac{(-1)^{n+1}\zeta(kn)}{n}=\ln 2+\sum_{n=1}^{\infty}\frac{(-1)^{n+1}(\zeta(kn)-1)}{n},$
where the $\ln 2$ arises from $\sum_{n=1}^{\infty}(-1)^{n+1}/n=\ln 2$. Hence $B(k)=\ln f(k)-\ln 2$.
For (28), separate AA into odd and even indices to get $A=h+E$. Since $B$ alternates signs, B=hEB=h-E. The remaining identities follow; in particular, h=AEh=A-E recovers Theorem 1. ∎
@@ 5 Connection to the Riemann zeta function
Separating the $m=0$ term in Proposition 9:
$h(k)=\bigl[\zeta(k)-1\bigr]+\underbrace{\sum_{m=1}^{\infty}\frac{\zeta\bigl(k(2m+1)\bigr)-1}{2m+1}}_{C(k)},$
@@ 6.3 Proof of the gamma formula
Integrating $C^{\prime}(k)$ from 11 to $k$ using Lemmas 15 and 16:
@@ 9.3 Uniqueness of $k=3$
Among all integers $k\geq 2$:
  • (a)
    For $k=3$, $f(k)/(2g(k))=3/2\in\mathbb{Q}$.
  • (b)
    For $k=2$, $f(k)/(2g(k))=\sinh(\pi)/\pi$ is transcendental.
  • (c)
    For odd $k\geq 5$, the cyclotomic telescoping mechanism that produces rationality at $k=3$ does not apply: specifically, $\Phi_{2k}(n+1)\neq\Phi_{k}(n)$ for $k\geq 5$ odd.
In particular, $k=3$ is the unique integer $k\geq 2$ for which $f(k)/(2g(k))$ is known to be rational.
@@
Part (a) is Theorem 29.
Part (b). For even $k\geq 2$, the product formulas of Theorem 5 express $f(k)$ and $g(k)$ in terms of $\sin$ and $\sinh$ evaluated at algebraic multiples of $\pi$. In particular, for $k=2$ we have $f(2)/(2g(2))=\sinh(\pi)/\pi=(e^{\pi}-e^{-\pi})/(2\pi)$. By Nesterenko’s theorem [5], $\pi$ and $e^{\pi}$ are algebraically independent over $\mathbb{Q}$, so any nonzero element of the field $\mathbb{Q}(\pi,e^{\pi})$ is transcendental. Since $\sinh(\pi)/\pi\in\mathbb{Q}(\pi,e^{\pi})\setminus\{0\}$, the quantity $f(2)/(2g(2))$ is transcendental, and in particular irrational. The argument for general even $k$ is analogous: the product $f(k)/(2g(k))$ lies in a field generated over $\mathbb{Q}$ by $\pi$ and exponentials $e^{\pi\alpha}$ for algebraic $\alpha$, and is expected to be transcendental; at minimum, it is not known to be rational in these cases.
Part (c). The telescoping that occurs for $k=3$ depends on the cyclotomic identity $\Phi_{6}(n+1)=\Phi_{3}(n)$, which is special to the quadratic case: $(n+1)^{2}-(n+1)+1=n^{2}+n+1$. For odd primes $p\geq 5$, the polynomials $\Phi_{p}$ and $\Phi_{2p}$ both have degree $\varphi(p)=p-1\geq 4$, and $\Phi_{2p}(n+1)\neq\Phi_{p}(n)$; for example, the coefficient of $n^{p-2}$ in $\Phi_{2p}(n+1)$ is $\binom{p-1}{1}-1=p-2$, whereas in $\Phi_{p}(n)$ it is $1$. Hence
@@
For even $k\geq 4$, the product $f(k)/(2g(k))$ lies in a field generated over $\mathbb{Q}$ by $\pi$ and values $e^{\pi\alpha}$ for algebraic $\alpha$, and is expected to be transcendental. However, Nesterenko’s theorem [5] establishes algebraic independence only for $\{\pi,e^{\pi},\Gamma(1/4)\}$, and extending the argument to all even $k$ would require algebraic independence results for $\{e^{\pi j/k}:1\leq j<k\}$ that are not currently available unconditionally.
Remark 34.
The transcendence of $f(k)/(2g(k))$ (established for $k=2$) does not imply the transcendence of $h(k)=\frac{1}{2}\ln\bigl(f(k)/(2g(k))\bigr)$: Baker’s theorem applies only when the argument of the logarithm is algebraic, and $\ln(\alpha)$ can be algebraic for transcendental $\alpha$ (e.g., $\ln(e^{\sqrt{2}})=\sqrt{2}$). Thus the arithmetic nature of $h(k)$ for even $k$ remains open. For odd $k\geq 5$, even the algebraic nature of $f(k)/(2g(k))$ is unresolved; should this quantity prove to be algebraic and not equal to $1$, Baker’s theorem would immediately yield the transcendence of $h(k)$.
@@ 10.5 OEIS sequence data
The decimal expansion of $\ln(\sinh(\pi)/\pi)=2h(2)$ appears in the OEIS as A393668. The Euler–Mascheroni constant $\gamma$ is A001620. The values of $\zeta(k)$ for small $k$ correspond to well-known constants: $\zeta(2)=\pi^{2}/6$ (A013661), $\zeta(3)$ is Apéry’s constant (A002117), $\zeta(4)=\pi^{4}/90$ (A013662), and so on.